OpenStudy (anonymous):

What is the remainder of this long division problem? 3x-2 then long hand with 6x^3+2x^2+11x-10 in it.

OpenStudy (amistre64):

when does 3x-2 = 0?

OpenStudy (amistre64):

if P(x) = (3x-2) Q(x) + R(x) then 3x-2 = 0, P(x) = R(x)

OpenStudy (anonymous):

So it would be a 0 remainder?

OpenStudy (amistre64):

dunno, when does 3x-2 = 0?

OpenStudy (amistre64):

since it is stated that: 6x^3+2x^2+11x-10 = (3x-2) Q(x) + R(x) then the value of x that makes 3x-2 = 0, lets say its u, will zero out the Q(x) part and we are left with 6u^3+2u^2+11u-10 = 0 + R(u)

OpenStudy (amistre64):

3x-2 = 0 when x=2/3 so the remainder has to be: 6(2/3)^3+2(2/3)^2+11(2/3)-10 = R(2/3)

OpenStudy (amistre64):

or just work thru the longhand division stuff ....